# Flow of Reasoning: Training LLMs for Divergent Problem Solving with Minimal Examples

### v4

@@ Abstract
The ability to generate diverse solutions to a given problem is a hallmark of human creativity. This divergent reasoning is also crucial for machines, enhancing their robustness and enabling them to assist humans in many applications such as scientific discovery. However, existing approaches to multi-step reasoning with large language models (LLMs) have mostly focused only on reasoning accuracy, without further discovering more diverse valid solutions. For example, supervised fine-tuning can improve LLM reasoning quality, but requires extensive supervised data to capture the full range of possible solutions. Reward-maximization reinforcement learning aims to find limited highest-reward solutions while neglecting the solution diversity. To fill this gap, we propose Flow of Reasoning (FoR), an efficient diversity-seeking LLM finetuning method aimed at improving reasoning quality and diversity with minimal data. FoR formulates multi-step LLM reasoning as a Markovian flow on a DAG-structured reasoning graph. This formulation allows us to incorporate and adapt principled GFlowNet approaches, for finetuning LLMs to sample divergent paths with probabilities proportional to the (unnormalized) reward of target problems. Extensive experiments show that, with limited training examples (e.g., 15 examples), FoR enables the discovery of diverse, creative, high-quality solutions, greatly outperforming a wide range of existing inference and training methods across six challenging reasoning tasks, including BlocksWorld (embodied reasoning), Game24 (math puzzle solving), Rubik’s Cube (spatial reasoning), 1D-ARC (abstraction reasoning), GSM8k (math reasoning), and ProntoQA (logical reasoning). Code is available at https://github.com/Yu-Fangxu/FoR.
@@ 1 Introduction
State-of-the-art reasoning with large language models (LLMs), however, has largely focused on improving only the problem-solving accuracy with the topmost solution, without moving a step further to discover more diverse valid solutions. Specifically, inference methods, such as CoT (chain of thought, Wei et al., 2022), ToT (Yao et al., 2024), RAP (Hao et al., 2023), and others (Chen et al., 2024b; Besta et al., 2024), rely heavily on the underlying pretrained LLM’s capability and decoding algorithms to obtain divergent reasoning solutions. Moreover, the search-based inference (Yao et al., 2024; Hao et al., 2023; Chen et al., 2024b; Besta et al., 2024) can be computationally costly when searching for multiple reasoning paths. On the other hand, finetuning methods improve the inherent abilities of the underlying LLMs. However, the popular supervised finetuning (SFT) (Yue et al., 2023; Yu et al., 2023c) often demands extensive supervision data to capture the full diversity of solutions, which can be costly to label in many applications. Alternatively, reinforcement learning (RL), such as proximal policy optimization (PPO, Schulman et al., 2017), trains LLMs to generate the highest-reward reasoning solution and overlooks solution diversity. As shown in the case study in Figure 1, limited solutions are found by the methods mentioned above.
@@
To overcome the limitations, we introduce Flow of Reasoning (FoR), a data-efficient approach that finetunes LLMs for divergent reasoning with only minimal data. FoR draws inspirations from generative flow networks (GFlowNets) for amortized diverse sampling (Bengio et al., 2021) that have been studied in different domains like molecule synthesis (Koziarski et al., 2024; Kim et al., 2024a) and operation scheduling (Zhang et al., 2023a). In particular, FoR enables diversity-seeking finetuning of multi-step LLM reasoning, to sample high-quality reasoning paths with probabilities proportional to the reward of target problems (as opposed to reward maximization in conventional RL). To this end, we formulate multi-step LLM reasoning from a Markovian flow perspective (Figure 2), where each reasoning step corresponds to an edge (action) that leads to the next node (state) in a flow graph. The reasoning process thus forms a flow that travels step-by-step from an initial state to the terminal states of the target problem. Based on this new formulation, we introduce the trajectory balance objective and adapt efficient exploration methods from the recent GFlowNet studies, enabling effective finetuning of LLMs to align with the task reward using only 15 input examples.
@@
FoR differs crucially from the recent GFlowNet applications on autoregressive sequence generation with or without LLMs (Hu et al., 2023a; Malkin et al., 2022a). In particular, contrary to the token-level modeling in the previous work, FoR introduces higher-level modeling at the granularity of reasoning steps. This combines the best of the GFlowNet sequence generation (Hu et al., 2023a; Malkin et al., 2022a) and the aforementioned search-based LLM reasoning (Yao et al., 2024; Hao et al., 2023) while overcoming their limitations, by enabling more flexible DAG-structured reasoning graphs, more efficient handling of complex multi-step reasoning problems, and thereby greatly improved reasoning quality and diversity as shown in §4.
@@
We evaluate the divergent problem-solving capability of the proposed approach on six puzzle-solving problems that have proven challenging for LLM reasoning, including BlocksWorld that involves embodied reasoning (Kambhampati et al., 2024), Game24 involving math puzzle solving (Yao et al., 2024), Rubik’s Cube involving spatial reasoning (Ding et al., 2023), 1D-ARC involving abstraction reasoning (Xu et al., 2023b), GSM8k involving math reasoning (Cobbe et al., 2021) and PrOntoQA involving logical reasoning (Saparov & He, 2022). Empirical results show that FoR, with limited (e.g. about 15) training examples, generates diverse, high-quality solutions, greatly outperforming a wide range of baselines with 20% - 85% improvements, including supervised training methods like SFT, reward-maximizing reinforcement learning like PPO, diversity-seeking approaches GFN-CoT and various decoding methods, and advanced inference methods like CoT, ToT, GoT, and RAP. Ablation studies further validate the key designs in FoR that lead to robustness and effectiveness.
@@ 2 Related Work
Divergent Thinking. Divergent thinking involves two branches: lateral and vertical (Waks, 1997). Lateral thinking explores unconventional ideas and challenges assumptions, as seen in brain teasers (Zhong et al., 2024; Summers-Stay et al., 2023), while vertical thinking focuses on structured, step-by-step reasoning. Our work focuses on vertical thinking, aiming to generate multiple different reasoning paths to solve a single problem.
LLM reasoning. Recent LLMs (Achiam et al., 2023; Touvron et al., 2023; Chowdhery et al., 2023) have shown strong potential in tackling complex reasoning tasks (Hu et al., 2023c; Zhang et al., 2023d; Yu et al., 2023b). (1) Fine-tuning LLMs, including supervised fine-tuning (SFT) and reinforcement learning (RL), is a key method for improving LLM reasoning abilities. SFT, leveraging large and high-quality datasets of reasoning chains, has proven highly effective (Yu et al., 2023c; Yue et al., 2023; Yuan et al., 2024a). RL techniques like PPO are widely used for optimizing reward-driven behavior in LLMs (Ouyang et al., 2022; Bai et al., 2022; Havrilla et al., 2024). However, both approaches tend to limit solution diversity. (2) Prompting-based methods engages LLMs in a step-by-step thinking process. Chain-of-Thought (CoT) (Wei et al., 2022) enhances LLM performance by guiding them through intermediate steps to reach the final answer. Building on CoT, methods like ToT (Yao et al., 2024) and GoT (Besta et al., 2024) model reasoning as tree and graph searches, enabling exploration of multiple paths. Other methods, like RAP (Hao et al., 2023) and XoT (Ding et al., 2023) use planning approaches such as MCTS to refine reasoning trajectories.
GFlowNets. GFlowNets (Bengio et al., 2021) were developed to generate diverse, high-reward samples from unnormalized distributions (Shen et al., 2023b; Roy et al., 2023; Zhang et al., 2023c; Ma et al., 2024; Pan et al., 2023b), making them particularly effective in domains like molecule synthesis (Koziarski et al., 2024; Kim et al., 2024a; Lu et al., 2024) and biological sequence design  (Ghari et al., 2023; Jain et al., 2022), where diversity is essential. Unlike traditional reinforcement learning (e.g., PPO), which focuses on maximizing reward, GFlowNets sample complete trajectories with probabilities proportional to their rewards, promoting exploration of the solution space. Recently, GFlowNets with LLMs have been applied to autoregressive tasks like token-level text generation  (Hu et al., 2023a), but these approaches are limited to token-level sampling, making them less suited for complex reasoning. FoR extends GFlowNet principles to higher-level multi-step reasoning, modeling it as a Markovian flow through a DAG, enabling the exploration of divergent reasoning paths.
@@
Figure 2: Left: The forward policy $P_{F}(s_{t}|s_{t-1};\theta,g)$ in the flow-based formulation is parameterized as LLM and finetuned with the trajectory balance objective (Eq.5) to achieve the desired flow $F(s_{n})=R(s_{n})$ on all terminal states $s_{n}$. Right: FoR incorporates local search with a destroy-and-reconstruction process to augment informative trajectories in training (§3.2.2). This facilitates efficient exploration and improves policy learning.
@@ 3.1 Multi-step LLM Reasoning as Markovian Flow
We first formulate the divergent multi-step reasoning tasks from a Markovian flow perspective (section 3.1). Then, we develop an efficient LLM training method using GFlowNet objectives for multi-step reasoning tasks (section  3.2). Multi-Step Reasoning Problem as DAG. Consider a multi-step reasoning problem that gives an initial state $s_{0}$ and a goal $g$. For example, in BlocksWorld (Figure 2), an initial state is the starting configuration of the block stack, and a goal describes the desired configuration of blocks. Reasoning aims to find complete paths (or trajectories) that lead from the initial state to the states that satisfy the goal. Given a current state $s$, applying an action on it leads to the transition to the next state $s^{\prime}$, denoted as $s\to s^{\prime}$. For example, in Figure 2, state $s_{0}$ transits to $s_{1}$ after an action "pickup blue". A complete trajectory is thus a sequence of transitions $\tau=(s_{0}\to s_{1}\to\dots\to s_{n})\in\mathcal{T}$, where $s_{n}$ is the terminal state and $\mathcal{T}$ is the set of all complete trajectories. Given a current state $s_{t}$, there could be multiple alternative next actions, resulting in different branches of the reasoning. Also, different sequences of actions can lead to the same intermediate/terminal states, as shown in Figure 1. As a result, the multi-step reasoning has the structure of a directed acyclic graph (DAG).
@@
The reasoning graph consists of diverse trajectories that lead to different terminal states. A crucial component often provided in reasoning tasks is the reward $R(s_{n})\in\mathbb{R}_{\geq 0}$, which assigns a numerical value to any terminal state $s_{n}$. For instance, a terminal state meeting the goal $g$ receives a high reward. As discussed in §1, to generate diverse high-quality reasoning trajectories for solving a task, we want to sample the trajectories with probabilities proportional to the reward. This significantly differs from popular reinforcement learning methods (e.g., PPO) and prompting-based planning algorithms (e.g., RAP, ToT), which focus on optimizing for only the maximum-reward trajectory.
The Flow Perspective.
@@
Sampling complex multi-step trajectories from the (often unnormalized) reward is particularly challenging (LeCun et al., 2006; Qin et al., 2022). To overcome the difficulty, we consider the above reasoning problem from a flow-based viewpoint which was initially developed in (Bengio et al., 2021) and has been studied in other machine learning settings like molecule generation (Pan et al., 2022; Malkin et al., 2022a; Shen et al., 2023a; Li et al., 2023c; Lahlou et al., 2023; Li et al., 2024a; He et al., 2024). Specifically, we define a trajectory flow function $F:\mathcal{T}\to\mathbb{R}_{\geq 0}$. Analogous to the classical concept of flows in networks, the flow $F(\tau)$ can be thought of as the volume of water traveling along this path $\tau$. Based on this, for any state $s$, we can define the state flow $F(s)=\sum_{s\in\tau}F(\tau)$, and for any edge $s\to s^{\prime}$, the edge flow $F(s\to s^{\prime})=\sum\nolimits_{s\to s^{\prime}\in\tau}F(\tau)$. These concepts of (unnormalized) flow are connected to the (normalized) probability distributions. Specifically, the flow trajectory determines a distribution over trajectories:
$\begin{split}P(\tau)=F(\tau)/Z,~{}~{}~{}Z=\sum\nolimits_{\tau\in\mathcal{T}}F(\tau).\end{split}$ (1)
@@
$\begin{split}P(\tau)=\prod\nolimits_{t=1}^{n}P_{F}(s_{t}|s_{t-1}),~{}~{}~{}\text{where}~{}P_{F}(s_{t}|s_{t-1})=F(s_{t-1}\to s_{t})/F(s_{t-1}).\end{split}$ (2)
That is, intuitively, $P_{F}(s_{t}|s_{t-1})$ characterizes the proportion of water at node $s_{t-1}$ that travels toward node $s_{t}$. The distribution $P_{F}$ is also called the forward policy, which can be used to generate a trajectory $\tau$ by sampling a sequence of transitions step-by-step starting from the initial state $s_{0}$. Equivalently (Bengio et al., 2023), there exists a backward policy that defines the distributions $P_{B}(\cdot|s_{t})$ over the parents of each state $s_{t}$: $P_{B}(s_{t-1}|s_{t})=F(s_{t-1}\to s_{t})/F(s_{t})$.
@@
$\begin{split}F(s_{n})=R(s_{n}),~{}~{}~{}\forall~{}\text{terminal state}~{}s_{n}.\end{split}$ (3)
@@
LLM Parameterization.
We parameterize the forward policy $P_{F}$ with an LLM and finetune as described in the next section. Specifically, for a reasoning task, we express its goal $g$, action $a$, and state $s$ as natural language (see Figure 1, BlocksWorld as an example). At each reasoning step $t$, the LLM generates an action $a_{t}\sim P_{\text{LLM}}(a|s_{t};\theta,g,c)$, where $c$ is an appropriate prompt (e.g., instructions or in-context demonstrations). The prompts used in the experiments are detailed in Appendix C. Once an action is generated, the state transits to the next $s_{t+1}=T(s_{t},a_{t})$ with a transition function $T$. Therefore, assuming that different actions applying to the same state $s_{t}$ lead to different next states, and that action $a_{t}$ leads to state $s_{t+1}$, we can write $P_{F}(s_{t+1}|s_{t};\theta,g)=P_{\text{LLM}}(a_{t}|s_{t};\theta,g,c)$. In the experiments, we follow previous work and define $T$ either by an LLM with appropriate prompts and greedy decoding (e.g., BlocksWorld as in (Hao et al., 2023)) or by the environment (e.g., Rubik’s Cube as in (Ding et al., 2023)).

### v6

# Flow of Reasoning: Training LLMs for Divergent Reasoning with Minimal Examples

@@ Abstract
The ability to generate diverse solutions to a given problem is a hallmark of human creativity. This divergent reasoning is also crucial for machines, enhancing their robustness and enabling them to assist humans in many applications such as scientific discovery. However, existing approaches to multi-step reasoning with large language models (LLMs) have mostly focused only on reasoning accuracy, without further discovering more diverse valid solutions. For example, supervised fine-tuning improves reasoning quality but requires vast labeled data, while reward-maximizing reinforcement learning finds top-reward solutions while neglecting the solution diversity. To fill this gap, we propose Flow of Reasoning (FoR), an efficient diversity-seeking LLM finetuning method aimed at improving reasoning quality and diversity with minimal data. FoR formulates multi-step LLM reasoning as a Markovian flow on a DAG-structured reasoning graph. This formulation allows us to incorporate and adapt principled GFlowNet approaches, for finetuning LLMs to sample divergent paths with probabilities proportional to the (unnormalized) reward of target problems. Extensive experiments show that, with limited training examples (e.g., 15 examples), FoR enables the discovery of diverse, creative, high-quality solutions, greatly outperforming a wide range of existing inference and training methods across six challenging reasoning tasks, including BlocksWorld (embodied reasoning), Game24 (math puzzle solving), Rubik’s Cube (spatial reasoning), 1D-ARC (abstraction reasoning), GSM8k (math reasoning), and ProntoQA (logical reasoning). Code is available at https://github.com/Yu-Fangxu/FoR.
Machine Learning, ICML
@@ 1 Introduction
State-of-the-art reasoning with large language models (LLMs), however, has largely focused on improving only the reasoning accuracy with the topmost solution, without moving a step further to discover more diverse valid solutions. Specifically, inference methods, such as CoT (chain of thought, Wei et al., 2022), ToT (Yao et al., 2024), RAP (Hao et al., 2023), and others (Chen et al., 2024b; Besta et al., 2024), rely heavily on the underlying pretrained LLM’s capability and decoding algorithms to obtain diverse reasoning solutions. Moreover, the search-based inference (Yao et al., 2024; Hao et al., 2023) can be computationally costly when searching for multiple reasoning paths. On the other hand, the finetuning methods like supervised finetuning (SFT) (Yue et al., 2023; Yu et al., 2023c) often demands extensive supervision data to capture the full diversity of solutions, which can be costly to label in many applications. Alternatively, reward-maximization reinforcement learning (RL), such as proximal policy optimization (PPO, Schulman et al., 2017), trains LLMs to generate the highest-reward reasoning solution and overlooks solution diversity. As shown in the case study in Figure 1, the above-mentioned methods generate limited diverse solutions.
@@
To overcome the limitations, we introduce Flow of Reasoning (FoR), a data-efficient approach that finetunes LLMs for diverse reasoning with minimal training data. FoR draws inspirations from generative flow networks (GFlowNets) for amortized diverse sampling (Bengio et al., 2021) that have been studied in different domains like molecule synthesis (Koziarski et al., 2024) and operation scheduling (Zhang et al., 2023c). In particular, FoR enables diversity-seeking finetuning of multi-step LLM reasoning, to sample high-quality reasoning paths with probabilities proportional to the reward of target problems (as opposed to reward maximization in conventional RL like PPO). To this end, we formulate multi-step LLM reasoning from a Markovian flow perspective (Figure 2), where each reasoning step corresponds to an edge (action) that leads to the next node (state) in a flow graph. The reasoning process thus forms a flow that travels step-by-step from an initial state to the terminal states of the target problem. Based on this new formulation, we use the trajectory-balanced objective and adapt efficient exploration methods from the recent GFlowNet studies, enabling finetuning of LLMs to generate more accurate and diverse solutions in reasoning tasks.
@@
FoR differs crucially from the recent GFlowNet applications on autoregressive sequence generation with or without LLMs (Hu et al., 2023a; Malkin et al., 2022a). In particular, contrary to the token-level modeling in the previous work (Hu et al., 2023a), FoR introduces higher-level modeling at the granularity of reasoning steps. This overcomes the limitations of aforementioned search-based LLM reasoning (Yao et al., 2024; Hao et al., 2023) by amortizing inference into training. Additionally, GFlowNets’ exploratory nature (Madan et al., 2024) enhances both reasoning quality and diversity, addressing the limitations of fine-tuning methods like SFT and PPO (see §4).
@@
We evaluate the divergent reasoning capability of our approach across six challenging benchmarks: BlocksWorld (embodied reasoning; Kambhampati et al., 2024), Game24 (math puzzles; Yao et al., 2024), Rubik’s Cube (spatial reasoning; Ding et al., 2023), 1D-ARC (abstraction; Xu et al., 2023b), GSM8K (math reasoning; Cobbe et al., 2021), and ProntoQA (logical reasoning; Saparov & He, 2022; see Appendix 4.7). Empirical results show that FoR, with limited (e.g. 15) training examples, generates diverse, high-quality solutions, greatly outperforming a wide range of baselines with 20% - 85% improvements, including supervised training methods like SFT, reward-maximizing RL like PPO, diversity-seeking approaches GFN-CoT and various decoding methods, and advanced inference methods like CoT, ToT, GoT, and RAP. Ablation studies further validate the key designs in FoR that lead to robustness and effectiveness.
@@ 2 Related Work
Figure 2: Left: The forward policy $P_{F}(s_{t}|s_{t-1};\theta,g)$ in the flow-based formulation is parameterized as LLM and finetuned with the trajectory balance objective (Eq.5) to achieve the desired flow $F(s_{n})=R(s_{n})$ on all terminal states $s_{n}$. Right: FoR incorporates local search with a destroy-and-reconstruction process to augment informative trajectories in training (§3.2). This facilitates efficient exploration and improves policy learning.
Divergent Thinking. Divergent thinking involves two branches: lateral and vertical (Waks, 1997b). Lateral thinking explores unconventional ideas and challenges assumptions, as seen in brain teasers (Zhong et al., 2024; Summers-Stay et al., 2023), while vertical thinking focuses on structured, step-by-step reasoning. Our work focuses on vertical thinking, aiming to generate multiple different reasoning paths to solve a single problem.
LLM Reasoning. Recent LLMs (Achiam et al., 2023) have shown strong potential in tackling complex reasoning tasks (Hu et al., 2023c; Zhang et al., 2023d; Yu et al., 2023b). (1) Fine-tuning LLMs, including supervised fine-tuning (SFT) and reward-maximization reinforcement learning (RL), is a key method for improving LLM reasoning abilities. SFT, leveraging large and high-quality datasets of reasoning chains, has proven highly effective (Yu et al., 2023c). RL techniques like PPO and DPO are widely used for optimizing reward-driven behavior in LLMs (Ouyang et al., 2022; Bai et al., 2022). However, both approaches tend to limit solution diversity. (2) Prompting-based methods engage LLMs in a step-by-step thinking process. Chain-of-Thought (CoT) (Wei et al., 2022) enhances LLMs performance by guiding them through intermediate steps to reach the final answer. Building on CoT, the methods like ToT (Yao et al., 2024) and GoT (Besta et al., 2024) model reasoning as tree and graph searches, and RAP (Hao et al., 2023) and XoT (Ding et al., 2023) use planning approaches to refine reasoning trajectories. DoT (Naik et al., 2023) enhances reasoning diversity through varied prompts.
GFlowNets. GFlowNets (Bengio et al., 2021) were developed to generate diverse, high-reward samples from unnormalized distributions (Shen et al., 2023b; Roy et al., 2023; Zhang et al., 2023b), making them particularly effective in domains like molecule synthesis (Koziarski et al., 2024; Kim et al., 2024a; Lu et al., 2024), where diversity is essential. Recently, GFlowNets have been integrated with LLMs (Hu et al., 2023a; Kwon et al., 2024; Lee et al., 2024b; Song et al., 2024), but none of these have addressed multi-step reasoning problems. While Takase et al. (2024) is a concurrent work, its scope is limited to math problems and does not extend to general reasoning problems. In contrast, FoR extends GFlowNets to multi-step reasoning, modeling it as a Markovian flow through a DAG, enabling the exploration of divergent reasoning paths across six reasoning benchmarks.
@@ 3.1 Multi-step LLM Reasoning as Markovian Flow
Multi-Step Reasoning Problem as DAG. Consider a multi-step reasoning problem that gives an initial state $s_{0}$ and a goal $g$. For example, in BlocksWorld (Figure 2), an initial state is the starting configuration of the block stack, and a goal describes the desired configuration of blocks. Reasoning aims to find complete paths (or trajectories) that lead from the initial state to the state(s) that satisfy the goal. Given a current state $s$, applying an action on it leads to the transition to the next state $s^{\prime}$, denoted as $s\to s^{\prime}$. For example, in Figure 2, the state $s_{0}$ transits to $s_{1}$ after an action "pickup blue". A complete trajectory is thus a sequence of transitions $\tau=(s_{0}\to s_{1}\to\dots\to s_{n})\in\mathcal{T}$, where $s_{n}$ is the terminal state and $\mathcal{T}$ is the set of all complete trajectories. Given a current state $s_{t}$, there could be multiple alternative next actions, resulting in different branches of the reasoning steps. Also, different sequences of actions can lead to the same intermediate/terminal states, as shown in Figure 1. As a result, we first formulate the multi-step reasoning problem as a directed acyclic graph (DAG).
@@
The reasoning graph consists of diverse trajectories that lead to different terminal states. A crucial component often provided in reasoning tasks is the reward $R(s_{n})\in\mathbb{R}_{\geq 0}$, which assigns a numerical value to any terminal state $s_{n}$. For instance, a terminal state meeting the goal $g$ receives a high reward. As discussed in §1, to generate diverse correct reasoning trajectories for solving a task, we want to sample the trajectories with probabilities proportional to the reward. This significantly differs from popular reinforcement learning methods (e.g., PPO) and prompting-based planning algorithms (e.g., RAP, ToT), which focus on optimizing for only the maximum-reward trajectory.
@@
The Flow Perspective. Sampling complex multi-step trajectories from the (often unnormalized) reward is particularly challenging (LeCun et al., 2006; Qin et al., 2022). To overcome the difficulty, we consider the above reasoning problem from a flow-based viewpoint which was initially developed in (Bengio et al., 2021) and has been studied in other machine learning settings like molecule generation (Li et al., 2023c; Lahlou et al., 2023; Li et al., 2024a; He et al., 2024). Specifically, we define a trajectory flow function $F:\mathcal{T}\to\mathbb{R}_{\geq 0}$. Analogous to the classical concept of flows in networks, the flow $F(\tau)$ can be thought of as the volume of water traveling along this path $\tau$. Based on this, for any state $s$, we can define the state flow $F(s)=\sum_{s\in\tau}F(\tau)$, and for any edge $s\to s^{\prime}$, the edge flow $F(s\to s^{\prime})=\sum\nolimits_{s\to s^{\prime}\in\tau}F(\tau)$. These concepts of (unnormalized) flow are connected to the (normalized) probability distributions. Specifically, the flow trajectory determines a distribution over trajectories:
$P(\tau)=F(\tau)/Z,~{}~{}~{}Z=\sum\nolimits_{\tau\in\mathcal{T}}F(\tau).$ (1)
@@
$P(\tau)=\prod\nolimits_{t=1}^{n}P_{F}(s_{t}|s_{t-1}),$ (2)
$\text{where}~{}P_{F}(s_{t}|s_{t-1})=F(s_{t-1}\to s_{t})/F(s_{t-1}).$
That is, intuitively, $P_{F}(s_{t}|s_{t-1})$ characterizes the proportion of water at node $s_{t-1}$ that travels toward node $s_{t}$. The distribution $P_{F}$ is also called the forward policy, which can generate a trajectory $\tau$ by sampling a sequence of transitions step-by-step starting from the initial state $s_{0}$. Equivalently (Bengio et al., 2023), there exists a backward policy that defines the distributions $P_{B}(\cdot|s_{t})$ over the parents of each state $s_{t}$: $P_{B}(s_{t-1}|s_{t})=F(s_{t-1}\to s_{t})/F(s_{t})$
@@
$\small\begin{split}F(s_{n})=R(s_{n}),~{}~{}~{}\forall~{}\text{terminal state}~{}s_{n}.\end{split}$ (3)
@@
LLM Parameterization. We parameterize the forward policy $P_{F}$ with an LLM and finetune as described in the next section. Specifically, for a reasoning task, we express its goal $g$, action $a$, and state $s$ as natural language (see Figure 1, BlocksWorld as an example). At each reasoning step $t$, the LLM generates an action $a_{t}\sim P_{\text{LLM}}(a|s_{t};\theta,g,c)$, where $c$ is an appropriate prompt (e.g., instructions or in-context demonstrations). The prompts used in the experiments are detailed in Appendix C. Once an action is generated, the state transits to the next $s_{t+1}=T(s_{t},a_{t})$ with a transition function $T$. Therefore, assuming that different actions applying to the same state $s_{t}$ lead to different next states, and that action $a_{t}$ leads to state $s_{t+1}$, we can write $P_{F}(s_{t+1}|s_{t};\theta,g)=P_{\text{LLM}}(a_{t}|s_{t};\theta,g,c)$. In the experiments, we follow previous work and define $T$ either by an LLM with appropriate prompts and greedy decoding (e.g., BlocksWorld as in (Hao et al., 2023)) or by the environment (e.g., Rubik’s Cube as in (Ding et al., 2023)).